59 A voltage has been applied across a capacitor. If the dielectric is replaced with another dielectric constant ight times as great and the voltage is reduced to half of what it was, the ENERGY STORED in the capacitor is how many times the original stored energy? A) 1/2 C) 1/4 D) 8 4

may you explain how you would find the answer of 2
Answer #1

energy stored in capacitor initially

E1 = 0.5 (k1 C) V1^2 ...(i)

now, energy stored later on

E2 = 0.5* k2 C V2^2 ... (ii)

given that

k2 = 8 k1

V2 = V1 / 2

putting in (ii)

E2 = 0.5* (8 k1) C V1^2/ 4

E2 = 2 ( 0.5* ( k C) V1^2)

from (i)

E2 = 2 E1

E2/E1 = 2

========

Comment in case any doubt, will reply for sure. Goodluck